What volume of oxygen will be left unreacted when a volume of 100cm3 of hydrogen and 200cm3 of oxygen are exploded in a eudiometer?
1
Expert's answer
2020-07-07T14:35:26-0400

The reaction between hydrogen and oxygen can be calculated as following:

H2 +O2 = H2O

First, the number of moles of the reactants must be calculated. As 1 mol of any gas occupies 22.4 L:

n(oxygen) = V(oxygen) / 22.4 L/mol = 200 cm3 / 22.4 L/mol = 0.2 L / 22.4 L/mol = 0.009 mol

n(hydrogen) = V(hydrogen) / 22.4 L/mol = 100 cm3 / 22.4 L/mol = 0.1 L / 22.4 L/mol = 0.0045 mol

As a result, hydrogen is a limiting reactant. Therefore, during the explosion, 0.0045 mol of hydrogen and 0.0045 / 2 = 0.00225 mol of oxygen will be used. From here, 0.009 mol - 0.00225 mol = 0.00675 mol of oxygen will be left.

Finally, the volume of oxygen left is:

V(left oxygen) = 0.00675 mol × 22.4 L/mol = 0.15 L = 150 cm3


Answer: 150 cm3